# PSR J1922+37: a 1.9-second pulsar discovered in the direction of the old open cluster NGC 6791

### v1

@@ Abstract
More than 300 pulsars have been discovered in Galactic globular clusters; however, none have been found in open clusters. Here we present results from 20-hour pulsar searching observations in seven open clusters with the Five-hundred-meter Aperture Spherical radio Telescope (FAST). Our first discovery is a 1.9-second pulsar (J1922+37) found in the direction of the old open cluster NGC 6791. The measured dispersion measure (DM) implies a distance of 4.79 kpc and 8.92 kpc based on the NE2001 and YMW16 electron density models, respectively. Given the large uncertainty of DM distance estimates, it is likely that PSR J1922+37 is indeed a member of NGC 6791, for which the distance is $4.19\pm 0.02$ kpc based on Gaia Data Release 3. If confirmed, PSR J1922+37 will be the first pulsar found in Galactic open clusters. We outline future observations that can confirm this pulsar-open cluster association and discuss the general prospects of finding pulsars in open clusters.
@@ 1 Introduction
Pulsars are well known for their fast spin, high density, and strong gravitational and magnetic field, rendering them astrophysical laboratories unachievable on Earth (see e.g. Lorimer & Kramer, 2012). Globular clusters have the highest spatial concentrations of pulsars, with more than 300 pulsars discovered in 45 globular clusters [^1] https://www3.mpifr-bonn.mpg.de/staff/pfreire/GCpsr.html. In particular, 47 Tucanae alone has more than 40 pulsars, while Terzan 5 has nearly 50 pulsars. The large number of pulsars has provided a unique tool to probe the density profile and the dynamics of their host clusters, and to constrain the mass of potential intermediate massive black holes inside (Freire et al., 2017; Prager et al., 2017; Perera et al., 2017).
The high yield of pulsars in globular clusters is believed to be caused by their high stellar encounter rate, which is directly related to their crowded stellar environments (Hui et al., 2010; Bahramian et al., 2013). Specifically, the encounter rate, $\Gamma$, is proportional to $\rho_{\rm c}^{2}r_{\rm c}^{3}/v$, where $\rho_{\rm c}$ and $r_{\rm c}$ are the density and size of the core, respectively, and $v$ is the velocity dispersion of the cluster (Verbunt, 2002; Verbunt & Freire, 2014), while the number of pulsar discoveries, $\lambda$, is expected to follow the scaling relation $\log\lambda=-1.1+1.5\log_{10}\Gamma$ (Turk & Lorimer, 2013). More recently, Yin et al. (2024) showed that escape velocity could also be a determining factor for $\lambda$.
@@
Open clusters also have a relatively high stellar density, and more than 2900 such clusters have been identified, about 10 times that of globular clusters (e.g. Bica et al., 2019, Table 1). However, no pulsar discoveries have been confirmed in any open clusters yet. The difference is probably caused by the lower stellar density and thus a lower encounter rate in open clusters. In addition, the lower escape velocity of open clusters also makes it more difficult to retain pulsars that usually suffer from natal kicks (e.g. Pfahl et al., 2002; Ivanova et al., 2008; Boyles et al., 2011). The short lifetime ($\sim 10^{8}$ yr) of open clusters also means that they are likely to have disintegrated before any pulsars were discovered inside. As a result, open clusters were unfavorable places for pulsar searching and were usually overlooked in past search efforts. However, in the search for unusual types of pulsar, open clusters deserve more attention. Any confirmed pulsar discovery in these clusters will greatly advance our understanding of stellar evolution and pulsar formation.
@@
As the largest single-dish telescope, the Five-hundred-meter Aperture Spherical radio Telescope (FAST) (Nan et al., 2011) provides an unprecedented sensitivity for pulsar searches. In addition, the adoption of a 19-beam receiver also provides a decent speed to observe a large sky area (Jiang et al., 2020). Thus, FAST has been proven to be a productive telescope for pulsar discovery. To date, FAST has discovered more than 800 new pulsars in the past few years, mainly through the Galatic Plane Pulsar Survey (GPPS, Han et al. 2021), the Commensal Radio Astronomy FAST Survey (CRAFTS, Li et al. 2018) and dedicated globular cluster searches (e.g. Pan et al., 2021).
@@
In this paper, we present results from a 20-h FAST observing project (ID: PT2021_0102), with the goal of finding exotic pulsar systems in seven open clusters. After initial analysis of these observations, we found a 1.9-second pulsar, J1922+37, which is likely to be associated with open cluster NGC 6791. If confirmed, J1922+37 will be the first pulsar discovered in an open cluster and will provide new insights into the stellar evolution in open clusters.
@@ 2 Observations
Using the Milky Way star cluster catalog [^2] https://heasarc.gsfc.nasa.gov/W3Browse/star-catalog/mwsc.html (Kharchenko et al., 2013), we chose seven open clusters that are old and massive as targets. The basic information of the clusters is shown in Table 1. We see that all clusters are within $\sim 10^{\circ}$ of the Galactic plane. If no pulsars are found in the open clusters, we may still be able to detect pulsars in the foreground or background sky; note that the same sky locations (of these open clusters) have been or will be covered by the GPPS survey (Han et al., 2021), which motivates us to adopt a longer integration time for our observations (see Tables 2) and 3 for details).
Observations were carried out using all beams of the L-band 19-beam receiver (Jiang et al., 2020) with a time resolution of 49.152 $\mu$s. Two polarizations and 4096 frequency channels were used throughout the observations. The receiver provides a bandwidth of 500 MHz, but only 400 MHz are effective. Data were recorded with the pulsar backend in search mode and saved in psrfits format (Hotan et al., 2004) with a default depth of 8 bits. For calibration purposes, each observation began with a low-intensity noise injection, which has a modulating period of 0.2 second and lasted for one minute.
Taking into account the gaps between individual beams, four pointings are necessary to cover a given region (e.g. Jiang et al., 2020). For efficient coverage, all clusters were observed using two groups of on-off mode observations [^3] https://fast.bao.ac.cn/cms/article/24/, except for IC 4756, which was observed with the more efficient snapshot mode, the same observation mode used for GPPS (Han et al., 2021). Due to the large angular size of IC 4756, two groups of snapshot-mode observations were applied to cover the central region of the cluster. The observation position and integration length of the clusters are given in Table 2 (for the on-off mode) and Table 3 (for the snapshot mode). For the on-off mode, we provide the central beam (M01) of each pointing, while for the snapshot mode, only the central beam of the first pointing is provided. In both modes, the position of the remaining beams can be deduced by accounting for the geometry of the 19-beam receiver and considering the slewing track of the receiver in each mode (Jiang et al., 2020; Han et al., 2021).
@@ 3 Search method
A bottleneck of processing pulsar search mode data is the slow I/O speed due to the large data volume. To reduce data size, we convert the data from 8 bits to 2 bits and halve the number of frequency channels using the digifil routine in the dspsr software package (van Straten & Bailes, 2011). The conversion thus reduces the size of the data by a factor of 8, but only leads to a few percent decrease in signal-noise-ratio (S/N), as demonstrated using the known pulsar J1837+0528g [^4] http://zmtt.bao.ac.cn/GPPS/GPPSnewPSR.html in the line-of-sight of IC 4756 (see Section 4.1). After conversion, the data have 2048 frequency channels while the time resolution remains unchanged.
To minimize the impact of radio frequency interference (RFI), we found the bad frequency channels in each observation using the rfifind routine in prestosoftware package (Ransom, 2011). A length of one second was used to integrate for the analysis. The bad frequency channels were masked out for the remaining process.
We also removed the periodic interference using the sigpyproc software package [^5] https://sigpyproc3.readthedocs.io/en/latest/. For this purpose, we first removed the bad frequency channels from each observation. For each beam, a time series was then generated at 0 DM and a Fast Fourier Transform (FFT) was performed to obtain the power spectrum using sigpyproc. Since the receiver’s 19 beams observe the sky simultaneously, any periodic interference would appear across multiple beams. Therefore, to identify such periodic radio frequency interference or “birdies”, we flagged frequency bins if they exceeded a certain threshold ($\sigma>4$) in the power spectrum of four or more beams. These frequencies were used as a common birdie list for all 19-beams of a single pointing.
@@
Both dedispersion and acceleration search were performed with the peasoup [^6] https://github.com/ewanbarr/peasoup software package, which can also take into account the bad frequency channels and birdies mentioned above. More importantly, peasoup is based on the Graphic Processing Unit (GPU), which can significantly speed up the search process (e.g. Morello et al., 2019; Sengar et al., 2023). For each search, in spectral domain, we set an S/N threshold of 6.5 and only kept the candidates above the threshold. As a result, 15000 candidates were found for each pointing.
To reduce the number of candidates to a manageable scale, we adopted a parameter-based method, which uses a combination of the candidate period, S/N and the number of harmonics to pick out the most probable pulsar candidates (Sengar et al., 2023). We then folded the search mode data using the period, DM, and acceleration, with the dspsr routine in the psrchive [^7] https://psrchive.sourceforge.net/ software package (Hotan et al., 2004). The resulting archive files were first cleaned with the interference removal program clfd [^8] https://github.com/itachi-gf/clfd/tree/master (Morello et al., 2023), before being processed with the pdmp routine in psrchive, which optimizes the period, DM, and S/N and provides diagnostic plots for the inspection. Using the refined parameters, the sifting and folding process was repeated once, further reducing the number of candidates and creating diagnostic plots for these candidates, which were visually inspected to pick out pulsars and highly confident candidates.
@@ 4 Results
4.1 Discovery and confirmation
Two pulsars were found in the search, PSRs J1837+05 and J1922+37. PSR J1837+05 is a 6.26 ms pulsar and has a DM of $\sim 120.8$ pc cm-3, consistent with PSR J1837+0528g, which was discovered by GPPS. J1837+05 was detected in three adjacent beams, which are also consistent with the discovery position of J1837+0528g, confirming the re-detection nature. J1837+05 is in the line-of-sight of open cluster IC 4756, but its DM is about 20 times the DM estimate using the cluster distance and the electron density models (see Table 1). J1837+05 is thus more likely in the background than in the open cluster.
PSR J1922+37 has a period of 1.92 s and a DM of $\sim 85$ pc cm-3 (Fig. 1), and is in the line-of-sight of open cluster NGC 6791 (Fig. 2). The pulsar was discovered on 22 January 2022 in two adjacent beams (M08) with a comparable S/N $\sim 60$. In addition to the fundamental spin period, harmonics at the 2nd, 3rd, 6.5th, 17th, 19th, 21st and 25th were also detected. No significant acceleration was detected. Using FAST, the pulsar was later confirmed in an 18-min follow-up observation and in a 6-min grid observation (see Section 4.3 for more details).
@@ 4.1 Discovery and confirmation
Figure 2: The position of J1922+37 and the member stars of NGC 6791. The filled circles in orange, pink, cyan, and olive show the tiled beams of all four observation pointings. The two discovery beams (M08) are indicated by the hollow stars, while the optimized position is labeled by the solid red star, and the cluster members are marked by the gray stars. The blue and red circles show the boundary of the tidal radius and Jacobi radius, respectively.
@@ 4.2 Pulse width and flux density
where we use $T_{\rm sys}=18$ K, $\Delta f=400$ MHz (Jiang et al., 2019), and $G$ is the effective gain. For the discovery observation, which has the highest S/N = 66, $n_{\rm p}=2$ and $t_{\rm int}=3400$ s. According to Jiang et al. (2020), the gain of beam M08 is about 0.92 times that of the central beam at 1250 MHz, which gives $G_{0}=14.8$ K Jy-1. Note that the center of the discovery beams is about half a beam-width off the source, so the effective gain should be smaller. Assume a Gaussian beam pattern (e.g. Condon & Ransom, 2016), then $G=G_{0}/2$. Using the parameters above, we obtain a flux density of $\sim 7~{}\mu$Jy.
All the observations mentioned above were spread over 3 years and thus are insufficient to produce a phase-coherent timing solution. As a result, spin-down and characteristic age are unavailable at the moment. No reliable polarization information was obtained, as all the observations either do not have sufficiently high S/N or full Stokes parameters, making the polarization measurement insignificant.
@@ 4.3 Position and distance
We refined the pulsar position using two groups of on-off-mode observations with FAST and the same receiver. The observations were centered on the midpoint of the two adjacent discovery beams and fully covered the nearby region in grids. Each pointing was 6 minutes. The pulsar was detected in three adjacent beams, one beam reporting an S/N of $\sim 17$ and two beams of $\sim 7$. Using the S/N-weighted sum of the center of the beams, we update the pulsar position to RA = 19:22:03 and Dec = +37:46:13 with an uncertainty of the beam size ($\sim 3\arcmin$).
According to the position, PSR J1922$+$37 is in the same direction as NGC 6791 with an offset of $\sim 14\arcmin$ from the center of the cluster. Fig. 2 shows the member stars of NGC 6791, which are taken from Hunt & Reffert (2024) and have a membership probability greater than 30% [^9] https://cdsarc.cds.unistra.fr/viz-bin/cat/J/A+A/686/A42#/article. PSR J1922+37 is slightly outside the King (1962) tidal radius of NGC 6791 but still within its Jacobi radius, an alternative to estimate the tidal radius (Binney & Tremaine, 2008; Hunt & Reffert, 2024).
Moreover, using the optimal DM and position, the NE2001 electron density model (Cordes & Lazio, 2002) gives a distance estimate of $d_{\rm NE2001}\sim 4.79$ kpc, which is close to the cluster distance reported by the Milky Way star cluster survey (4.93 kpc, Kharchenko et al. 2013) and that inferred from Gaia Data Release 2 ($4.45\pm 0.10$ kpc, Dias et al., 2021), but slightly differs from that obtained using red clump giants in the cluster ($4.02\pm 0.15$ kpc Gao & Chen, 2012) and that using Gaia Data Release 3 ($4.19\pm 0.02$ kpc, Hunt & Reffert, 2024). However, the distance estimated by the YMW16 electron density model (Yao et al., 2017) is $d_{\rm YMW16}\sim 8.92$ kpc, about two times the distance to NGC 6791.

### v2

@@ Abstract
More than 300 pulsars have been discovered in Galactic globular clusters; however, none have been found in open clusters. Here we present results from a 20-hour survey of seven open clusters with the Five-hundred-meter Aperture Spherical radio Telescope (FAST). Our first discovery is a 1.9-second pulsar (J1922+37) found in the direction of the old open cluster NGC 6791. The measured dispersion measure (DM) implies a distance of 4.79 kpc and 8.92 kpc based on the NE2001 and YMW16 electron density models, respectively. Given the large uncertainty of DM distance estimates, it is plausible that PSR J1922+37 is indeed a member of NGC 6791, for which the distance is $4.19\pm 0.02$ kpc based on Gaia Data Release 3. If confirmed, PSR J1922+37 will be the first pulsar found in a Galactic open cluster. We outline future observations that can confirm this pulsar-open cluster association and discuss the importance of this pulsar for calibrating the characteristic ages of the pulsar population.
@@ 1 Introduction
Pulsars are well known for their fast spin, high density, and strong gravitational and magnetic field, rendering them astrophysical laboratories unachievable on Earth (see e.g. Lorimer & Kramer, 2012). Globular clusters have the highest spatial concentrations of pulsars, with more than 300 pulsars discovered in 45 globular clusters [^1] https://www3.mpifr-bonn.mpg.de/staff/pfreire/GCpsr.html. In particular, 47 Tucanae alone has more than 40 pulsars, while Terzan 5 has nearly 50 pulsars. The large number of pulsars has provided a unique tool to probe the density profile and the dynamics of their host clusters, and to constrain the mass of potential intermediate massive black holes inside (Freire et al., 2017; Prager et al., 2017; Perera et al., 2017). The high yield of pulsars in globular clusters is believed to be caused by their high stellar encounter rate, which is directly related to their crowded stellar environments (Verbunt, 2002; Hui et al., 2010; Bahramian et al., 2013; Verbunt & Freire, 2014).
@@
Open clusters also have a relatively high stellar density, and more than 2900 such clusters have been identified, about 10 times the number of globular clusters (e.g. Bica et al., 2019, and their Table 1). However, no radio pulsars have been discovered in any open clusters yet [^2] An X-ray pulsar and magnetar, CXO J164710.2--455216, was discovered in the Galactic young massive cluster Westerlund 1 (Muno et al., 2006). Based on epicyclic approximation and simulations, PSR 1932+1059 was proposed to be a runaway system from the open cluster IC 4665 (Bobylev, 2008).. The difference is probably due to their smaller cluster mass, which is usually several orders of magnitude smaller than that of globular clusters. Considering the total number of open clusters and globular clusters and assuming a similar formation rate of neutron star per unit mass, this means fewer neutron stars could be present in open clusters. In addition, the lower escape velocity of open clusters also makes it more difficult to retain pulsars that usually suffer from natal kicks (e.g. Pfahl et al., 2002; Ivanova et al., 2008; Boyles et al., 2011). The short lifetime ($\sim 10^{8}$ yr) of open clusters also means that they are likely to have disintegrated before any pulsars were discovered inside. As a result, open clusters were unfavorable places for pulsar searching and were usually overlooked in past search efforts. However, in the search for unusual types of pulsar, open clusters deserve more attention. Any confirmed pulsar discovery in these clusters will greatly advance our understanding of stellar evolution and pulsar formation. It would tell us that some pulsars can be born with very small kicks, and also permit a comparison of the pulsar’s characteristic age with that of the open cluster.
@@
As the largest single-dish telescope, the Five-hundred-meter Aperture Spherical radio Telescope (FAST) (Nan et al., 2011) provides an unprecedented sensitivity for pulsar searches. In addition, the adoption of a 19-beam receiver also provides a decent survey speed to observe a large area (Jiang et al., 2020). Thus, FAST has been proven to be a productive telescope for pulsar discovery. To date, FAST has discovered more than 1000 new pulsars in the past few years, mainly through the Galactic Plane Pulsar Survey (GPPS, Han et al. 2021), the Commensal Radio Astronomy FAST Survey (CRAFTS, Li et al. 2018) and dedicated globular cluster searches (e.g. Pan et al., 2021).
@@
In this paper, we present results from a 20-h FAST observing project (ID: PT2021_0102), with the goal of finding exotic pulsar systems in seven open clusters. After initial analysis of these observations, we found a 1.9-second pulsar, J1922+37, which may be associated with the open cluster NGC 6791. If confirmed, J1922+37 will be the first pulsar discovered in an open cluster and will provide new insights into stellar evolution and neutron star production in open clusters.
@@ 2 Observations
Using the Milky Way star cluster catalog [^3] https://heasarc.gsfc.nasa.gov/W3Browse/star-catalog/mwsc.html (Kharchenko et al., 2013), we chose seven open clusters that are old and massive as targets. The basic information of the clusters is shown in Table 1. We see that all clusters are within $\sim 10^{\circ}$ of the Galactic plane. If no pulsars are found in the open clusters, we may still be able to detect pulsars in the foreground or background sky; note that the same sky locations (of these open clusters) have been or will be covered by the GPPS survey (Han et al., 2021), which motivated us to adopt a longer integration time for our observations (see Tables 2 and 3 for details).
Observations were carried out using all beams of the L-band 19-beam receiver (Jiang et al., 2020) with a time resolution of 49.152 $\mu$s. Two polarizations and 4096 frequency channels were used throughout the observations. The receiver provides a bandwidth of 500 MHz, but only 400 MHz are used in pulsar searches. Data were recorded with the pulsar backend in search mode and saved in psrfits format (Hotan et al., 2004) with a default bit depth of 8. For calibration purposes, each observation began with a low-intensity noise injection, which has a modulating period of 0.2 second and lasted for one minute.
Taking into account the gaps between individual beams, four pointings are necessary to cover a given region (e.g. Jiang et al., 2020). For efficient coverage, all clusters were observed using two groups of on-off mode observations [^4] https://fast.bao.ac.cn/cms/article/24/, except for IC 4756, which was observed with the more efficient snapshot mode, the same observation mode used for the GPPS survey (Han et al., 2021). Due to the large angular size of IC 4756, two groups of snapshot-mode observations were applied to cover the central region of the cluster. The observation position and integration length of the clusters are given in Tables 2 and 3 for the on-off mode and the snapshot mode, respectively. For the on-off mode, we provide the position of the central beam (M01) of each pointing, while for the snapshot mode, only the central beam of the first pointing is provided. In both modes, the position of the remaining beams can be deduced by accounting for the geometry of the 19-beam receiver and considering the slewing track of the receiver in each mode (Jiang et al., 2020; Han et al., 2021).
@@ 3 Search method
A bottleneck of processing pulsar search-mode data is the slow I/O speed due to the large data volumes. To reduce data size, we converted the data from 8 bits to 2 bits and halved the number of frequency channels using the digifil routine in the dspsr software package (van Straten & Bailes, 2011). The conversion thus reduced the size of the data by a factor of 8, but only leads to a few percent decrease in signal-noise-ratio (S/N), as demonstrated using the known pulsar J1837+0528g [^5] http://zmtt.bao.ac.cn/GPPS/GPPSnewPSR.html in the line-of-sight of IC 4756 (see Section 4.1). After conversion, the data have 2048 frequency channels while the time resolution remains unchanged.
To minimize the impact of radio frequency interference (RFI), we found the bad frequency channels in each observation using the rfifind routine in presto software package (Ransom, 2011). A length of one second was used to integrate for the rfifind analyses. The bad frequency channels were masked out for the remaining processing.
We also removed the periodic interference using the sigpyproc software package [^6] https://sigpyproc3.readthedocs.io/en/latest/. For this purpose, we first removed the bad frequency channels from each observation. For each beam, a time series was then generated at zero DM and a Fast Fourier Transform (FFT) was performed to obtain the power spectrum using sigpyproc. Since the receiver’s 19 beams observe the sky simultaneously, any periodic interference would appear across multiple beams. Therefore, to identify such periodic radio frequency interference or “birdies”, we flagged frequency bins if they exceeded a certain threshold ($\sigma>4$) in the power spectrum of four or more beams. These frequencies were used as a common birdie list for all 19-beams of a single pointing.
@@
Both de-dispersion and acceleration searches were performed with the peasoup [^7] https://github.com/ewanbarr/peasoup software package, which can also take into account the bad frequency channels and birdies mentioned above. More importantly, peasoup is based on the Graphic Processing Unit (GPU), which can significantly speed up the search process (e.g. Morello et al., 2019; Sengar et al., 2023, 2024). For each search, in the spectral domain we set an S/N threshold of 6.5 and only kept the candidates above that threshold. As a result, about 15000 candidates were found for each pointing.
To reduce the number of candidates to a manageable scale, we adopted a parameter-based method, which uses a combination of the candidate period, S/N and the number of harmonics to pick out the most probable pulsar candidates as in Sengar et al. (2023). We then folded the remaining candidates’ search-mode data using the period, DM, and acceleration, with the dspsr routine in the psrchive [^8] https://psrchive.sourceforge.net/ software package (Hotan et al., 2004). The resulting archive files were first cleaned with the interference removal program clfd [^9] https://github.com/itachi-gf/clfd/tree/master (Morello et al., 2023), before being processed with the pdmp routine in psrchive, which optimizes the period, DM, and S/N and provides diagnostic plots for human inspection. Using the refined parameters, the sifting and folding process was repeated again, further reducing the number of candidates and creating diagnostic plots for these candidates, which were visually inspected to pick out any known pulsars and high significance candidates.
@@ 4 Results
4.1 Discovery
Two pulsars were detected in the search, PSRs J1837+05 and J1922+37. PSR J1837+05 is a known 6.26 ms pulsar and has a DM of $\sim 120.8$ pc cm-3, consistent with PSR J1837+0528g, which was discovered by GPPS. J1837+05 was detected in three adjacent beams, which are also consistent with the discovery position of J1837+0528g, confirming the association with the known pulsar. J1837+05 is in the line-of-sight of open cluster IC 4756, but its DM is about 20 times larger than the DM estimate using the cluster distance and the electron density models (see Table 1). J1837+05 is thus more likely in the background than in the open cluster.
PSR J1922+37 has a period of 1.92 s and a DM of $\sim 85$ pc cm-3 (Fig. 1), and is in the line-of-sight of open cluster NGC 6791 (Fig. 2). The pulsar was discovered on 22 January 2022 in two adjacent beams (M08) with a comparable S/N $\sim 60$. In addition to the fundamental spin period, many harmonics were also detected. No significant acceleration was detected. Using FAST, the pulsar was later confirmed in an 18-min follow-up observation and in a 6-min gridded observation (see Section 4.3 for more details).
@@ 4.1 Discovery
Figure 2: The position of J1922+37 and the member stars of NGC 6791. The filled circles in orange, pink, cyan, and olive show the tiled beams of all four observation pointings. The two discovery beams (M08) are indicated by the hollow stars, while the optimized position is labeled by the solid red star, and the cluster members are marked by the gray stars. The blue and red circles show the boundaries of the tidal and Jacobi radii respectively.
@@ 4.2 Pulse width and flux density
where we use $T_{\rm sys}=18$ K, $\Delta f=400$ MHz (Jiang et al., 2019), and $G$ is the effective gain. For the discovery observation, which has the highest S/N = 66, $n_{\rm p}=2$ and $t_{\rm int}=3400$ s. According to Jiang et al. (2020), the gain of beam M08 is about 0.92 times that of the central beam at 1250 MHz, which gives $G_{0}=14.8$ K Jy-1. Note that the center of the discovery beams is about half a beam-width off the source, so the effective gain should be smaller. If we assume a Gaussian beam pattern (e.g. Condon & Ransom, 2016), then $G=G_{0}/2$. Using the parameters above, we obtain a flux density of just $\sim 7~{}\mu$Jy.
The sparse number of observations mentioned above were spread over 3 years and are thus insufficient to produce a phase-coherent timing solution. As a result, spin-down and characteristic age are unavailable at the moment. No reliable polarization information was obtained, as all the observations either do not have sufficiently high S/N or full Stokes parameters, making the polarization measurement impossible.
@@ 4.3 Position and distance
We refined the pulsar position using two groups of on-off-mode observations with FAST and the same receiver. The observations were centered on the midpoint of the two adjacent discovery beams and fully covered the nearby region in grids. Each pointing was 6 minutes in duration. The pulsar was detected in three adjacent beams, one beam reporting an S/N of $\sim 17$ and two beams of $\sim 7$. Using the S/N-weighted sum of the center of the beams, we updated the pulsar position to RA = 19:22:03 and Dec = +37:46:13 with a conservative uncertainty of the beam diameter ($\sim 3\arcmin$).
According to the position, PSR J1922$+$37 is in the same direction as NGC 6791 with an offset of $\sim 14\arcmin$ from the center of the cluster. Fig. 2 shows the member stars of NGC 6791, which are taken from Hunt & Reffert (2024) and have a membership probability greater than 30% [^10] https://cdsarc.cds.unistra.fr/viz-bin/cat/J/A+A/686/A42#/article. NGC 6791 has a half-mass radius of $r_{\rm h}=4\farcm 42$ (Platais et al., 2011). PSR J1922+37 is slightly outside the King (1962) tidal radius ($r_{\rm t}=11\farcm 6$) of NGC 6791 but still within its Jacobi radius ($r_{\rm J}=29\farcm 7$), an alternative to estimate the radius of gravitational influence for cluster retention (Binney & Tremaine, 2008; Hunt & Reffert, 2024).
Using the optimal DM and position, the NE2001 electron density model (Cordes & Lazio, 2002) gives a distance estimate of $d_{\rm NE2001}=4.79\pm 0.96$ kpc (assuming a nomial 20% uncertainty, e.g. Straal et al. 2020), which is close to the cluster distance, $d_{\rm NGC6791}$, reported by the Milky Way star cluster survey (4.93 kpc, Kharchenko et al. 2013) and that inferred from Gaia Data Release 2 ($4.45\pm 0.10$ kpc, Dias et al., 2021), but slightly differs from that obtained using red clump giants in the cluster ($4.02\pm 0.15$ kpc Gao & Chen, 2012) and that using Gaia Data Release 3 ($4.19\pm 0.02$ kpc, Hunt & Reffert, 2024). However, the distance estimated by the YMW16 electron density model (Yao et al., 2017) is $d_{\rm YMW16}=8.9\pm 1.8$ kpc, about two times the distance to NGC 6791. Clearly we cannot definitively argue for the association based solely on the DM.